# CardioLive: Empowering Video Streaming with Online Cardiac Monitoring

### v1

@@ 5 Evaluation
Data Collection: There is no existing dataset that can fit our requirements, with audio-visual pairs and clear heart rate ground truth. Therefore, we self-collect the dataset through 8 commodity devices which span multiple mainstream platforms (iOS, Android, Windows, Mac), major brands, and device types (smartphones, tablets, laptops, and webcams) released between 2018-2024: Logitech C930 Webcam, OPPO Reno 2Z, Redmi Note 5, Honor 20i, MacBook Air M2, iPad 2018, iPad 2023 Pro, and iPhone 14. We leverage Polar H10 [8] to collect the ground truth. We utilize Flutter to develop a cross-platform application capable of recording raw video and audio, while also establishing the connections to the ground truth via MQTT. Our dataset comprises recordings from 10 users of diverse genders and regions. They are requested to read 10 materials [52], counting for 2,800 words. Each round lasts for 40 minutes. They wear the heartbeat belt when they are reading. We do not restrict users to a fixed distance from the recording devices. We leverage a tripod along with a ring light to cast different light sources on the users. We collected a total of 84,666 data clips, which are clipped into facial regions with 4-s windows. We resize the video frames to 72$\times$72$\times$3 and the audio is resampled to 8kHz. The missing frames will be duplicated adopting the same scheme as §4.3 mentions. We gain IRB from our university board. Moreover, we also make use of two publicly available video-only datasets: PURE [51] and MMPD [54] to validate the video-only solution.
@@
Software: We implement CardioNet through Pytorch 2.4.0. The model is trained via a single-card NVIDIA A100 80GB. We train the model with the learning rate of 1e-3, AdamW optimizer, batch size of 16, OneCycle scheduler. We use Pytorch JIT to compile the model. We write 2000+ lines C++ code to implement the service in Zoom and 1500+ lines JavaScript code for developing the service in the extension.
@@
Deployment: We propose two deployment paradigms, web-based and app-based. For the web-based one, we develop a browser extension that operates CardioLive in the background, which continuously captures audio and video data for processing, with results displayed on a canvas within the interface. In the app-based deployment, we register a bot in compliance with the policies of the video streaming companies. This bot joins the sessions as a member, similar to other participants, with the consent of all members. The data hook extracts audio and video towards inference engines. The processed results are delivered through a notification system. Notably, the inference can be performed either on the company’s cloud server or locally on the user’s device. In our real-world evaluation, we perform inference on the user’s device (Xiaoxin 16 Pro with AMD Ryzen 7 5800H), demonstrating the robustness and efficiency of our model.
5.1 Comparative Study
@@
We compare our CardioNet with various baselines. We choose the SOTA video-only baselines: TS-CAN [40], DeepPhys [21], PhysNet [72], EfficientPhys [41], RhythmFormer [77], POS [61]. The last one represents the signal processing based rPPG approaches. We also reimplement VocalHR [67], the recent work that employs human speech for detecting heart rate. Through this study, we will justify our superior performances using both audio and video modalities.
@@
Distances: We first experiment with different distances, ranging from 0.5m to 2.5m. We apply the logarithmic scale to each graph, with the base of 10. As shown in Fig. 13, while the error increases with distance for all methods, our approach consistently outperforms other baseline models at all tested distances. CardioNet achieves a MAE of just 1.40 BPM at 0.5m, significantly lower than the SOTA video-based baseline, i.e., RhythmFormer, by 73.7 %, and 96.7% lower than the worst-performing model, i.e., POS. Meanwhile, the audio-based model VocalHR has a MAE of 8.12 BPM at the same distance, which is 82.8% higher than ours. Even at the maximum testing distance of 2.5 meters, CardioNet is still 63.1% better than RhythmFormer and 77.9% better than VocalHR. This demonstrates the fusion of audio and video signals in CardioNet significantly enhances the overall performance. Besides, we observe the identical patterns of MAE, MAPE and RMSE, we will mainly report MAE for simplicity.
@@
Angles: We evaluate our model across a range of angles from 0° to $\pm$60° at a distance of 1 meter, as shown in Fig. 14. As the viewing angle increases, video-based methods suffer from significant performance degradation due to reduced visibility of facial features. However, CardioNet, through audio-visual fusion, maintains robust performance across all angles. While the video quality deteriorates with extreme angles, audio signals remain largely unaffected by viewing angles. Even at extreme angles like $\pm$60°, where video signals typically falter, our model achieves up to 38.9% lower MAE compared to baseline models, highlighting its superior resilience in challenging conditions. This result underscores the critical role of the audio modality in compensating for the loss of visual information at extreme angles.
@@
Noise Levels: We test heart rate estimation under noise levels from 30 dB to 38 dB. As in Fig. 15, increasing noise leads to higher absolute error. Nonetheless, CardioNet consistently outperforms the SOTA audio-only model VocalHR. This can be attributed to our temporal frequency filter design and the video modality which provides complementary information that remains stable under acoustic noise. For instance, at 30 dB, our model achieves a MAE of 1.25 BPM, significantly lower than VocalHR’s 8.64 BPM, and maintains this advantage even at 38 dB. The fusion network learns to adaptively reduce reliance on noisy audio features while leveraging more stable visual cues. The CDF curves show that CardioNet achieves higher cumulative probabilities at lower error thresholds, indicating its resilience to noise.
Noise Sources: We analyze the impact of noise sources such as rain, music, and TV shows in Fig. 16. CardioNet demonstrates strong noise resilience, particularly with rain noise, where it significantly outperforms VocalHR, achieving a MAE of just 1.94 BPM compared to 12.93 BPM. Even with more complex noise like music and TV shows, our model maintains lower MAEs, showcasing its robustness in diverse acoustic environments. This highlights the effectiveness of video modalities when facing the ambient noises.
Body Motions: Body motion can significantly impact the performance of heart rate detection models. To validate the robustness of our approach under different body motion scenarios, we evaluate the model in three typical body movements: walking, left-right (LR) rotation, and up-down (UD) rotation, as in Fig. 17. Despite the motion artifacts, CardioNet maintains robust performance, achieving an MAE of 1.35 BPM in the UD scenario, and consistently outperforms baselines by significant margins in all motion types. Our model benefits from the unique design of the motion-aware aggregation and temporal differentiation block. These prove the robustness of our model against body motions by effectively employing video plus audio modalities.
@@
5.2 Micro Benchmarks
Different Light Conditions: We assess our model under varying lightness levels from 0.3702 to 0.3259 in Fig.19(a), by adjusting the ring light. As lightness decreases, the MAE increases from 4.85 BPM to 8.16 BPM. This trend suggests that poorer conditions impact accuracy due to the reduced visibility of facial features. However, the model remains sufficiently robust, indicating that while lighting plays a role, the audio-visual fusion helps mitigate the negative effects.
@@ 5.2 Micro Benchmarks
Different FPS: We examine the model across various video frame rates (FPS), ranging from 30 FPS to 15 FPS, as shown in Fig.19(b). We interpolate the frame rate by adopting the principles discussed in §4.3. The model performs best at 30 FPS with an MAE of 1.75 BPM. Even at lower frame rates, particularly 15 FPS, the MAE increases to 4.56 BPM, while still remaining in the low level. This robust performance is achieved through our frame interpolation scheme and the audio branch’s ability to provide continuous cardiac information regardless of video frame rate. Also, our temporal differential block and irregular sampled time embedding block are equally vital to handle varying frame rates.
Different Quality of Image: We analyze the model’s performance under different video compression qualities, from 100 (highest quality) to 40 (lowest quality), as shown in Fig.19(c). Interestingly, the MAE does not consistently worsen with lower quality. At extreme compression levels, the model achieves the lowest MAE of 2.49 BPM, potentially due to smoothing effects that enhance key facial features. This suggests that while high compression degrades visual information, moderate to high levels of compression might benefit the model by reducing noise.
@@
(c) Quality of Image
Figure 19: The performances under different light, FPS and image conditions.
@@
Different Environments: Our model’s performance is evaluated across various environmental settings, including Office, Outdoor, Conference Room, and Laboratory, as shown in Fig.20(a). The model performs best in the Office environment with a MAE of 1.40 BPM. Notably, the latter three environments are not in the training set, yet the model maintains strong performance, demonstrating that our feature extraction generalizes well to unseen conditions.
Different Face Filters: We test various facial filters, including Smooth Face, Tint Skin, Adjust Brightness, Add Contrast, and Sharpen Face, as shown in Fig.20(b). The Tint Skin filter yields the best performance with an MAE of 2.38 BPM, while a more aggressive filter like Sharpen Face achieves an MAE of 8.69 BPM. It shows our model effectively handles appearance changes while maintaining accuracy.
@@
Multi-person Scenarios: We evaluate the multi-person scenarios to justify the effectiveness of our preprocessing. We set the maximum number of people to be separated as two and crop the face region to a size of 72$\times$72 pixels. In our test, two users read materials simultaneously while sitting next to each other. We apply the facial and sound separation and match their audio and face regions. The test results show an MAE of 7.83 BPM and 8.13 BPM for each person, respectively. Although we observe some performance drops, our method still effectively distinguishes between the two individuals. Notably, the heart rates of the two people vary over time, with average heart rates of 76.17 BPM and 68.55 BPM, respectively ,showing our system can track distinct physiological states simultaneously.
@@
5.3 CardioLive in the wild
In this section, we will evaluate how CardioLive works as a service. We assess the service on two ends: the meeting platform and the online content providers.
@@ 5.3 CardioLive in the wild
Meeting Platforms: We choose Zoom as one of online meeting platforms, which provides the external developers the SDK to acquire access to the raw data. The average FPS is 28.4. We exploit the data hooks to acquire the streams and leverage buffer queues to hold the packets, as described in §4.2. The model consumes on average in 850ms on CPU. We choose a step size of 1s, and a window size of 4s. It means every second, we feed the 4-s windows for inference. The overall system latency averages 1.03 seconds, as depicted in Fig. 23(b). Notably, latency was primarily elevated at the start due to the initial model warm-up period [37]. This means our systems can run inference in real-time. Furthermore, we calculate the throughput of the whole system. We measure the time since the last update of heart rate. As we are feeding 4-s window of video and audio frames, the throughput is calculated as the volume of video and audio data processed per update period. As in Fig. 23(b), the average throughput of the system is 115.97 FPS, which is prominently larger than the common video FPS. It means that our systems can hold the service robustly without any freezes.
Online Content Providers: Online content providers such as YouTube often host their services in the web browser. We implement such a service in a Chrome extension. We employ the data hook developed from WebGL and WebAudio to acquire the streams. The average FPS is 26.97. The overall latency of our service is 1.23s, comparable to our step size 1s, as can be observed from Fig. 23(a). Meanwhile, the average throughput is 98.16 FPS, with a maximum throughput of 114.41 FPS. These results also justify our service will run smoothly in the extensions.
6 Related Work

### v2

@@ IV-C Service Design
[…]
IV-D Preprocessing
In this section, we will discuss the preprocessing pipelines. We use the OpenCV face detector to find faces. We also perform voice activity detection to segment the talking period. Additionally, we need to separate multiple persons, if any, and match their audio and videos.
Audio-Visual Matching: To realize the matching between speaking clips and facial hints, we adopt a cross-attention scheme [tao2021someone, jiang2023target]. Specifically, after the encoders, we get two features $M_{a}$ and $M_{v}$. These features are expected to encapsulate relevant speaking activities by employing temporal encoders [hu2018squeeze, afouras2018conversation]. To fuse the audio and video features, the audio features $M_{a}$ are integrated with the video data by treating $M_{v}$ as the target for querying through an attention framework. Conversely, the video features $M_{v}$ interact with $Q_{a}$, representing the audio query sequences. The outputs are concatenated together along the temporal direction.
V Evaluation
In this section, we systematically evaluate CardioLive. We perform comparison studies with the state-of-the-art (SOTA) video-based solutions and audio-based solutions. We mainly leverage our self-collected dataset. We use the following metrics to evaluate the accuracy of the model: Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Mean Absolute Percentage Error (MAPE).
@@ V Evaluation
Data Collection: There is no existing dataset that can fit our requirements, with audio-visual pairs and clear heart rate ground truth. Specifically, BP4D+ [zhang2016multimodal] provides additional IR images but not the necessary audio; MMSE-HR offers only video; and while MAHNOB-HCI contains audio, it does not require participants to speak—only incidental utterances are present, which does not fit our scope. Therefore, we self-collect the dataset through 8 commodity mobile and laptop devices. We leverage Polar H10 [PolarH10Polar] to collect the ground truth. We recruit 10 users of diverse genders and skin colors [^1] We have gained IRB from our university board.. They are requested to read 10 materials [sun2021ultrase]. Each round lasts for 40 minutes. We use a tripod along with a ring light to cast different light sources on the users. We collect overall of 84,666 data clips, which are clipped into facial regions with 4-s windows. We resize the video frames to $\times72$$\times$3 and the audio is resampled to 8kHz. The missing frames will be duplicated adopting the same scheme as §IV-C mentions. Moreover, we also make use of two publicly available video-only datasets: PURE [stricker2014non] and MMPD [10340857].
@@
Software: We implement CardioNet through Pytorch. The model is trained via a single-card NVIDIA A100 80GB. We train the model with the learning rate of 1e-3, AdamW optimizer, batch size of 16, and OneCycle scheduler. We use JIT to compile the model. We write 2000+ lines C++ code to implement the service in Zoom and 1500+ lines of JavaScript code for developing the service in the extension.
@@
Deployment: We propose two deployment paradigms, web-based and app-based. For the web-based one, we develop a browser extension that operates CardioLive in the background, which continuously captures audio and video data for processing, with results displayed on a canvas within the interface. In the app-based deployment, we register a bot in compliance with the policies of the video streaming companies, which joins the sessions as a member, with the consent of all members. The data hook extracts audio and video for inference engines. The processed results are delivered through a notification system. Notably, the inference can be performed either on the company’s cloud server or locally on the user’s device. In our real-world evaluation, we perform inference on the end device to show the robustness and efficiency. A demo video of CardioLive can be found here: https://youtu.be/xoLmxPD264g.
V-A Comparative Study
@@
We compare our CardioNet with various baselines. We choose the SOTA video-only baselines: TS-CAN [liu2020multi], DeepPhys [chen2018deepphys], PhysNet [yu2019remote], EfficientPhys [liu2023efficientphys], RhythmFormer [zou2024rhythmformer], POS [wang2016algorithmic]. The last one is the signal processing method. We also reimplement VocalHR [xu2022hearing], the recent work that employs human speech for detecting heart rate. Through this study, we will justify our superior performances using both audio and video modalities.
@@
Distances: We first experiment with different distances from 0.5m to 2.5m. We apply the log-10 scale to each graph. As shown in Fig. 14, while the error increases with distance for all methods, our approach consistently outperforms other baseline models at all tested distances. CardioNet achieves a MAE of just 1.40 BPM at 0.5m, significantly lower than the SOTA video-based baseline, i.e., RhythmFormer, by 73.7 %, and 96.7% lower than the worst-performing model, i.e., POS. Meanwhile, the audio-based model VocalHR has an MAE of 8.12 BPM at the same distance, which is 82.8% higher than ours. Even at the maximum testing distance of 2.5 meters, CardioNet is still 63.1% better than RhythmFormer and 77.9% better than VocalHR. This demonstrates that the fusion of audio and video signals in CardioNet significantly enhances the overall performance. Besides, we observe the identical patterns of MAE, MAPE and RMSE, we will mainly report MAE for simplicity.
@@
Angles: We evaluate our model across a range of angles from 0° to $\pm$60° at 1 meter, as shown in Fig. 15. As it increases, video-based methods suffer from significant performance degradation due to reduced visibility of facial features. However, CardioNet, through audio-visual fusion, maintains robust performance across all angles. While the video quality deteriorates with extreme angles, audio signals remain unaffected by viewing angles. Even at $\pm$60°, where video signals typically falter, our model achieves up to 38.9% lower MAE compared to baseline models. This result underscores the critical role of the audio modality at extreme angles.
@@
Noise Levels: We test heart rate under noise levels from 30 dB to 38 dB. As in Fig. 16, increasing noise leads to higher error. Nonetheless, CardioNet consistently outperforms the SOTA audio-only model VocalHR. This can be attributed to our temporal frequency filter design and the video modality, which provides complementary information that remains stable under acoustic noise. For instance, at 30 dB, CardioLive achieves a MAE of 1.25 BPM, significantly lower than VocalHR’s 8.64 BPM, and maintains this advantage even at 38 dB. The fusion network learns to adaptively reduce reliance on noisy audio features while keeping stable visual cues. The CDF curves show that CardioNet achieves higher cumulative probabilities at lower error thresholds, indicating its resilience to noise.
Noise Sources: We analyze the impact of noise sources such as rain, music, and TV shows in Fig. 17. CardioNet demonstrates strong noise resilience, particularly with rain noise, where it significantly outperforms VocalHR, achieving a MAE of just 1.94 BPM compared to 12.93 BPM. Even with more complex noise like music and TV shows, our model maintains lower MAEs, showcasing its robustness in diverse acoustic environments. This highlights the effectiveness of video modalities when facing ambient noise.
Body Motions: Body motion can significantly impact the performance of heart rate detection models. To validate the robustness of our approach, we evaluate the model in three typical body movements: walking, left-right (LR) rotation, and up-down (UD) rotation, as in Fig. 18. Despite the motion artifacts, CardioNet maintains robust performance, achieving an MAE of 1.35 BPM in the UD scenario, and consistently outperforms baselines by significant margins in all motion types. Our model benefits from the unique design of the motion-aware aggregation and temporal differentiation block. These prove the robustness of our model against body motions by effectively employing video plus audio modalities.
@@
Cross-Dataset Performance: To validate the generalizability of the model, we perform cross-dataset experiments on the current SOTA video-only solution and ours. As shown in Tab. I, we find that our model is significantly better when training on PURE and testing on MMPD, with an MAE of 2.845 BPM. This is because MMPD is a complicated dataset, where the motion and the light varies a lot. This can also be seen from Fig. 4, which shows that when confronted with different motions and lights, previous methods fail to provide a robust way to handle them. Conversely, our method incorporates motion-aware and frequency-aware modeling, which will enhance the performance. Furthermore, to our knowledge, no previous work has reported results for training on MMPD and testing on PURE. Our method achieves an MAE of 3.675 BPM and an RMSE of 8.07 in this setting. These results demonstrate the generalizability of our approach across different scenarios.
TABLE I: Cross-Dataset Performances on PURE and MMPD datasets.
Model Train-Set Test-Set MAE RMSE
TS-CAN [liu2020multi] PURE MMPD 13.93 15.14
PhysNet [yu2019remote] PURE MMPD 13.93 15.61
PhysFormer [yu2023physformer++] PURE MMPD 14.57 16.73
DeepPhys [chen2018deepphys] PURE MMPD 16.92 18.54
EfficientPhys [liu2023efficientphys] PURE MMPD 14.03 15.31
RhythmFormer [zou2024rhythmformer] PURE MMPD 8.98 14.85
Ours PURE MMPD 2.845 6.688
Ours MMPD PURE 3.675 8.07
@@
TABLE II: Ablation study evaluating contributions of each component. “w/o” denotes ablative removal; “w/o raw audio” denotes replacing raw audio input with Mel-Spectrogram.
MAE RMSE MAPE
w/o audio 8.400 $\pm$ 0.284 8.898 $\pm$ 5.810 0.113 $\pm$ 0.004
w/o ITE 3.168 $\pm$ 0.388 5.028 $\pm$ 3.486 0.045 $\pm$ 3.486
w/o FCB 3.660 $\pm$ 0.456 5.421 $\pm$ 4.212 0.048 $\pm$ 0.006
w/o FCB + w/o ITE 4.353 $\pm$ 0.473 6.014 $\pm$ 4.888 0.063 $\pm$ 0.007
w/o Raw Audio 4.168 $\pm$ 0.221 5.914 $\pm$ 2.314 0.061 $\pm$ 0.003
Ours (w/ all) 1.746 $\pm$ 0.380 4.114 $\pm$ 4.516 0.024 $\pm$ 0.005
V-B Ablation Study
To validate the contributions of each component in the model, we perform a comprehensive ablation study as shown in Tab. II.
w/o Audio: We first evaluate the model’s performance without the audio modality. As shown in the results, the MAE increases sharply from 1.746 to 8.400, indicating a significant performance drop. This demonstrates that audio–video fusion is highly effective, and that our temporal attention fusion mechanism successfully leverages complementary cues from both modalities.
w/o Irregular Time Embedding (ITE): The Irregular Time Embedding (ITE) component is designed to address the irregular sampling inherent in real-world video data. Removing this component results in a performance deterioration of 81.4% in MAE, indicating that temporal embedding is essential for handling unconstrained video streams.
w/o Frequency-Aware Conv Block (FCB): The Frequency-Aware Conv Block (FCB) is specifically designed to enhance the model’s ability to capture subtle frequency information and to aggregate it with spatial representations. Excluding FCB leads to a 52.3% increase in MAE, confirming its positive impact on model accuracy.
w/o ITE + w/o FCB: When both ITE and FCB are simultaneously removed, model performance degrades further compared to removing either module alone. This suggests that ITE and FCB independently and jointly enhance the model’s ability to process complex signals.
w/o Raw Audio: As discussed in §III, we propose using raw audio instead of the Mel-Spectrogram based on empirical observations. In this experiment, we replace the raw audio input with features extracted from Mel-Spectrograms using a ResNet encoder. The MAE increases from 1.746 BPM (raw audio) to 4.168 BPM (Mel-Spectrogram), confirming that raw audio is a more effective input representation for our task. Nonetheless, the inclusion of Mel-Spectrogram features still yields better results than removing the audio modality entirely, further highlighting the importance of audio in video-based cardiac monitoring systems compared to video-only solutions.
V-C Micro Benchmarks
Different Light Conditions: We assess our model under varying lightness levels from 0.3702 to 0.3259 in Fig.20(a), by adjusting the ring light. As it decreases, the MAE increases from 4.85 BPM to 8.16 BPM. This trend suggests that poorer conditions impact accuracy due to the reduced visibility of facial features. However, the model remains sufficiently robust, indicating that while lighting plays a role, the audio-visual fusion helps mitigate the negative effects.
Different FPS: We examine the model across various video frame rates, ranging from 30 to 15 FPS, as shown in Fig.20(b). We interpolate the frame rate by adopting the principles discussed in §IV-C. The model performs best at 30 FPS with an MAE of 1.75 BPM. Even at lower frame rates, particularly 15 FPS, the MAE increases to 4.56 BPM, while still remaining in a low level. This performance is achieved through our frame interpolation scheme and the audio branch’s ability to provide continuous cardiac information. Also, our temporal differential block and irregularly sampled time embedding block are equally vital to handle varying frame rates.
Different Quality of Image: We analyze the performance under various video compression qualities, from 100 to 40 (lowest quality), as shown in Fig.20(c). The MAE does not consistently worsen with lower quality. At extreme compression levels, the model achieves the lowest MAE of 2.49 BPM, potentially due to smoothing effects that enhance key facial features. This suggests that while high compression degrades visual information, moderate to high levels of compression might benefit the model by reducing noise.
@@ V-C Micro Benchmarks
(c) QoI
Figure 20: The performances under different light, FPS and quality of image.
@@
Different Environments: Our model’s performance is evaluated across various environmental settings, including Office, Outdoor, Conference Room, and Laboratory, as shown in Fig.21(a). The model performs best in the Office environment with an MAE of 1.40 BPM. Notably, the latter three environments are not in the training set, yet the model maintains strong performance, demonstrating that our feature extraction generalizes well to unseen conditions.
Different Face Filters: We test various facial filters, including Smooth Face, Tint Skin, Adjust Brightness, Add Contrast, and Sharpen Face, as shown in Fig.21(b). The Tint Skin filter yields the best performance with an MAE of 2.38 BPM, while a more aggressive filter like Sharpen Face achieves an MAE of 8.69 BPM. It shows our model effectively handles appearance changes.
@@
Multi-person Scenarios: We evaluate the multi-person scenarios to justify the effectiveness of our preprocessing. We set the maximum number of people to be separated as two and crop the face region to a size of $\times72$ pixels. In our test, two users read materials simultaneously while sitting next to each other. We apply the facial and sound separation and match their audio and face regions. The test results show an MAE of 7.83 BPM and 8.13 BPM for each person, respectively. Although we observe some performance drops, our method still effectively distinguishes between the two individuals. Notably, the heart rates of the two people vary over time, with average heart rates of 76.17 BPM and 68.55 BPM, respectively, showing our system can track distinct physiological states simultaneously.
@@
V-D CardioLive in the wild
In this section, we will evaluate how CardioLive works as a service.
Meeting Platforms: We choose Zoom as one of the online meeting platforms, which provides the external developers with the SDK to acquire access to the raw data. The average FPS is 28.4. We exploit the data hooks to acquire the streams and leverage buffer queues to hold the packets, as described in §IV-B. The model consumes on average in 850ms on CPU with a step size of 1s and an inference window size of 4s. The overall system latency averages 1.03 seconds, as depicted in Fig. 24(b). Notably, latency was primarily elevated at the start due to the initial model warm-up period [lion2016don]. This means our systems can run inference in real-time. Furthermore, we calculate the throughput of the whole system. We measure the time since the last update of the heart rate. As we are feeding a 4-s window of video and audio frames, the throughput is calculated as the volume of video and audio data processed per update period. As in Fig. 24(b), the average throughput of the system is 115.97 FPS, which is prominently larger than the common video FPS. It means that our systems can hold the service robustly without any freezes.
@@ V-D CardioLive in the wild
Online Content Providers: Online content providers such as YouTube often host their services in the web browser. We implement such a service in a Chrome extension. We employ the data hook to acquire the streams. The average FPS is 26.97. The overall latency of our service is 1.23s, comparable to our step size 1s, as can be observed from Fig. 24(a). Meanwhile, the average throughput is 98.16 FPS, with a maximum throughput of 114.41 FPS. These results also justify our service will run smoothly in the extensions.
Model Size: Our model contains 81.58M parameters and requires 7.398G FLOPs, which are comparable to other audio-video learning frameworks [gong2022contrastive]. We have further pruned and quantized the model for faster inference. Note that we must wait for the first window before processing; however, this initial latency is standard and acceptable for this type of task in the literature.
VI Related Work